OpenStudy (anonymous):

The time required to finish a test in normally distributed with a mean of 40 minutes and a standard deviation of 8 minutes. What is the probability that a student chosen at random will finish the test in between 32 and 56 minutes? 16% 82% 34% 2% so this is normal distribution and i don't understand it at allll

OpenStudy (anonymous):

for some reason i can't get normal distribution

OpenStudy (anonymous):

you need to find z corresponding to 32 and 56 respectively z= (x- mean)/sd for 32 wehave z1 = (32 - 40)/8= -1 for 56 z2 = (56-40)/8 = 2

OpenStudy (anonymous):

now you need the area from the normal distribution table between z2 and z1

OpenStudy (anonymous):

oh oh ok!

OpenStudy (anonymous):

im understanding it. slowly. but its coming

OpenStudy (anonymous):

34?